# Koopman Operators in Robot Learning

### v1

@@ V-B1 Optimization-based Methods
    Enforcing subspace invariance under the Koopman operator: Minimizing the EDMD’s residual error in (34)-(35) does not necessarily lead to a close-to-invariant subspace. Figure 6 illustrates this point with an example of a non-invariant subspace whose residual error can be made arbitrarily small depending on the selected basis of functions. Therefore, optimization problems (34)-(35) might lead to models not suitable for long-term prediction.
@@
    The work in [71] addresses this problem by providing an alternative loss function that captures the quality of the subspace. The notion of temporal forward-backward consistency index (or consistency index for brevity) is defined as 999 This definition is equivalent but different from the definition in [71].
@@
    The consistency index has several important properties. (1) $\mathcal{I}_{C}\in[0,1]$; (2) Unlike the cost functions in (34)-(35), $\mathcal{I}_{C}$ only depends on the subspace $\operatorname{span}(\mathbf{\Psi})$ and not any specific basis (Fig. 7 illustrates this); (3) Under specific change of basis, $\mathcal{I}_{C}$ can be viewed as the maximum eigenvalue of a positive semidefinite matrix ($I-K_{F}K_{B}$ is not generally symmetric) enabling the use of common optimization solvers; (4) Importantly, it provides a tight upper bound on the relative prediction error of EDMD on the entire subspace. Formally,
@@
    Figure 6: EDMD’s residual error for linear system $x^{+}=0.6x$ and the family of dictionaries $D_{\beta}(x)=[x,x+\beta\sin(x)]$ with $\beta\in[0.01,100]$. Note that for all $\beta\in\mathbb{R}\setminus\{0\}$, all dictionaries span the same subspace $\mathcal{S}=\operatorname{span}\{x,\sin(x)\}$. The residual error depends on the choice of basis for subspace $\mathcal{S}$. More importantly, $\mathcal{S}$ is not Koopman-invariant but the residual error can be arbitrarily close to zero depending on the basis.
    Figure 7: The square root of consistency index for the system and family of dictionaries of Fig. 6. Unlike EDMD’s residual error depicted in Fig. 6, the consistency index does not depend on the choice of basis and accurately measures the approximation quality of the subspace.
@@
    which minimizes the maximum EDMD function prediction error over the subspace. Note that, by minimizing the consistency index, one accounts for accurate prediction of all uncountably many members of the function space, as opposed to only finitely many functions in the optimizations (34)-(35). Therefore, by minimizing the consistency index, one can expect a more accurate prediction, especially in the long term. Figure 8 illustrates this point in a nonlinear pendulum example.
@@
    Learning a Change of Coordinates: The previous optimization methods aim to learn the dynamics in the original coordinates. However, the most intuitive coordinate system is generally not the best. A better approach would be to choose Koopman eigenfunctions as the basis for our coordinates along which dynamical evolution is linear, i.e. we can choose the injective (so it can be inverted) vector-valued function $\bm{\Phi}(x):\mathbb{R}^{N_{x}}\to\mathbb{R}^{N_{\bm{\Phi}}}$ that spans an (approximately) invariant subspace as our new coordinates and build an inverse map $\bm{\Theta}:\mathcal{R}(\bm{\Phi})\to\mathbb{R}^{N_{x}}$, where $\mathcal{R}(\bm{\Phi})$ is the range of $\bm{\Phi}$ and ideally we have $\bm{\Theta}\big{(}\bm{\Phi}(x)\big{)}=x$ for all $x\in\mathbb{R}^{N_{x}}$. Therefore, the dynamics in the new coordinates can be written as
    $\bm{\Phi}(x^{+})=A\bm{\Phi}(x)\;,$
    $x^{+}=\bm{\Theta}\big{(}\bm{\Phi}(x^{+})\big{)}\;,$
    where $A\in\mathbb{R}^{{N_{\bm{\Phi}}}\times{N_{\bm{\Phi}}}}$. Note that if $\bm{\Phi}$ is injective, we can work in the $\bm{\Phi}$ coordinates and do not necessarily need to use the inverse map $\bm{\Theta}$. However, analytically checking whether a map is injective or not is computationally intensive or even intractable (see e.g., [72]); therefore, the maps $\bm{\Phi}$ and $\bm{\Theta}$ are often learned simultaneously from data as a part of an auto-encoder neural network structure [73, 74, 75]. The work in [76] formulates a similar problem using rank-constrained semi-definite programs and provides convex relaxations to identify the linear embedding and the mapping that leads to the system’s states.
    Stability of attractors and dictionary learning: An important system property is the stability of attractors. This phenomenon is directly related to the existence of certain Koopman eigenfunctions [77]. Even though the Koopman operator is infinite-dimensional, the stability of the attractors is generally decidable by finitely many eigenfunctions. Moreover, these eigenfunctions lend themselves to a Lyapunov formulation, thus enabling the use of already existing rich literature of Lyapunov-based tools. The work [78] studies the stability problem from the angle of contractions and shows the equivalence of contraction and Koopman-based results regarding stability. Using the connection of Koopman eigenfunctions to attractors’ stability, [79] provides a neural network-based method to learn Lyapunov functions from data. Moreover, [80] provides a numerical method to build Lyapunov functions using Koopman eigenfunctions leading to estimates for the regions of attractions and their geometric features. In the context of dictionary/subspace learning, by enforcing the stability in the learned model (given the knowledge about the attractor’s stability), one can ensure a more realistic and accurate model [81, 82, 83].
    V-B2 Algebraic Search for Koopman-Invariant Subspaces
@@
    The work [36] provides a data-driven necessary and almost surely sufficient condition to identify all Koopman eigenfunctions in an arbitrary finite-dimensional function space. This result directly relies on the eigendecomposition of forward and backward EDMD matrices (cf. (V-B1)). Moreover, [36] provides an efficient and provably correct algebraic algorithm termed Symmetric Subspace Decomposition (SSD) that finds the maximal Koopman-invariant subspace in arbitrary finite-dimensional function spaces (termed “search space”). In addition, [37] provides a parallel version of the SSD algorithm suitable for searching through high-dimensional spaces. Given as search space the finite-dimensional function space spanned by the dictionary $\mathbf{\Psi}_{s}$, any basis $\mathbf{\Psi}$ with elements in $\operatorname{span}(\mathbf{\Psi}_{s})$ can be described using a matrix $C$ with full column rank as $\mathbf{\Psi}^{T}(\cdot)=\mathbf{\Psi}_{s}^{T}(\cdot)C$. If the subspace $\operatorname{span}(\mathbf{\Psi})$ is invariant under the Koopman operator, this gets reflected in the data, $\mathcal{R}(\mathbf{\Psi}(X)^{T})=\mathcal{R}(\mathbf{\Psi}(Y)^{T})$ (where $\mathcal{R}(\cdot)$ stands for the range space of its argument), which can also be written as
@@
    V-C Controller Design
    V-C1 Model-based Controller
    A system model plays a pivotal role in facilitating the design of model-based controllers, enabling us to exploit model predictions for the selection of appropriate control inputs aligned with a given task. As elucidated in Section V, we can derive both linear and nonlinear models from measurements using Koopman operator theory, seamlessly integrating them into the architecture of model-based controllers. The centerpiece and critical determinant impacting the efficacy of these controllers is the inherent “model.” The Koopman operator, adept at accommodating new real-time measurements and/or pre-existing offline measurements, can serve as a defining model constraint within the formulation of the controller structure. The incorporation of the Koopman operator into model-based control was initially introduced in [84]. Subsequently, the concepts of Koopman Model Predictive Control (Koopman MPC) [43] and Koopman Nonlinear Model Predictive Control (Koopman NMPC) [29] ensued. Through the process of estimating and refining models, utilizing either online measurements or offline learning gleaned from data acquired within physical environments, a Koopman-based model controller can effectively accomplish control objectives even for systems with incomplete knowledge or that are entirely unknown.
    In a Koopman-based optimal controller, control inputs are determined by iteratively solving
    subject to ${\bm{z}}_{i+1}=F_{\mathcal{K}}\left({\bm{z}}_{i},{\bm{u}}_{i}\right)$
    ${\bm{z}}_{0}=\mathbf{\Psi}(\bm{x}_{t})\;,$ (40)
    where $N_{h}$ is the number of time steps in the controller horizon, $J$ is a quadratic cost function, $\mathbf{\Psi}$ is the lifting dictionary, $F_{\mathcal{K}}$ is the Koopman-based system model, and $\bm{z}_{i}$ and $\bm{u}_{i}$ are the lifted system state and input at the $i^{\text{th}}$ time-step in the horizon, respectively. The problem’s objective function penalizes deviations from a desired trajectory, and its constraints enforce consistency with a Koopman system model. With a linear Koopman model realization, the optimization (V-C1) becomes a quadratic program (QP)
    subject to ${\bm{z}}_{i+1}=K_{A}{\bm{z}}_{i}+B{\bm{u}}_{i}$
    ${\bm{z}}_{0}=\mathbf{\Psi}(\bm{x}_{t})\;.$ (41)
    As discussed in Section IV-A, the Koopman operator can be used to describe a system’s dynamics in either linear or nonlinear form. A major advantage of a linear Koopman model realization is that it makes the formulation to be convex. This implies it has a unique globally optimal solution that can be efficiently computed without initialization even for high-dimensional models [85, 86, 87], making Koopman MPC well-suited for real-time feedback control applications. Nonlinear/bilinear Koopman model realizations render (V-C1) non-convex, which are less efficient to solve and may only yield locally optimal solutions [88]. Sometimes a nonlinear Koopman realization generates more accurate predictions, in which case such a trade-off is warranted. Bilinear Koopman model realizations have been explored recently to combine some of the advantages of linear and nonlinear models [27].
@@ V-C1 Model-based Controller
    V-C2 Active Learning
    The linear structure of the Koopman operator yields several advantages for learning the dynamics of a robotic system. The closed-form solution of the linear least-squares model fitting problem (16), (17), and (19) to approximate the Koopman operator can be leveraged to formulate active-learning controllers [30]. Specifically, assume that the Koopman dynamics from the dictionary of observables $\mathbf{\Psi}(x_{t+1})=K_{A}\mathbf{\Psi}(x_{t})+B\mathbf{u}_{k}$ form the mean of a normally distributed state in the lifted space $p(\mathbf{z}_{t+1}|K,\mathbf{z}_{k})$. Then, one can form an approximation of the Fisher-Information Matrix
    $\mathbf{I}=\mathbb{E}\left[\frac{\partial}{\partial K}\log p(\mathbf{z}_{t+1}|K,\mathbf{z}_{k})\frac{\partial}{\partial K}\log p(\mathbf{z}_{t+1}|K,\mathbf{z}_{k})^{\top}\right].$ (42)
    $\mathbf{I}=\frac{\partial\mathbf{z}_{t+1}}{\partial K}^{\top}\Sigma^{-1}\frac{\partial\mathbf{z}_{t+1}}{\partial K}\leq\text{Var}[K^{\star}]^{-1}\;,$ (43)
    where $\mathbf{z}_{t+1}=K_{A}\mathbf{z}_{t}+B\mathbf{u}_{t}$, and $\text{Var}[K^{\star}]^{-1}$ is the posterior variance of the approximate Koopman operator (as calculated by the Grammian matrix in (17)). The Fisher-information is a matrix that lower-bounds the posterior uncertainty in estimation, defined as the Cramér-Rao bound [89, 90]. The Fisher-information Matrix is differentiable and actionable, i.e. one can optimize controllers that directly maximize the Fisher information, thus improving the best-case posterior variance on the Koopman operator, given the current operator.
    $\underset{\{\mathbf{u}_{i}\}_{i=0}^{N_{h}}}{\text{minimize}}$ $\sum_{i=0}^{N_{h}-1}\mathfrak{I}(\mathbf{z}_{i},{{}^{t}}K)+\mathbf{u}_{i}^{\top}R\mathbf{u}_{t}$
    subject to $\mathbf{z}_{i+1}={{}^{t}}K_{A}\mathbf{z}_{i}+{{}^{t}}B\mathbf{u}_{i}$
    $\mathbf{z}_{0}=\mathbf{\Psi}(x_{t})$ (44)
    where $R\succ 0$ is a positive definite matrix, $\mathfrak{I}(\mathbf{z}_{i},{{}^{t}}K)$ is an optimality condition (e.g., D-, or T- Optimality) that reduces the matrix into a scalar value, $N_{h}$ is the time-horizon, and the superscript t indicates the current estimate of the Koopman operator given state-control data collected in the past. Note that the above optimization problem is done in a receding-horizon manner to account for changes in the Koopman operator. Controllers that are formulated from the optimization (V-C2) effectively lower the overall posterior variance of the model yielding an effective Koopman operator model with a few data points. The resulting model can be used across different tasks such as for an aerial robot to quickly recover from an unstable tumble or a legged robot to learn complex interaction models with granular media (see [30]).
    The simple structure of the Koopman operator has advanced learning in other directions. In particular, the use of deep models to approximate the function observables has made significant strides in expanding the use-case of methods based on the Koopman operator [30, 68, 91]. Deep function observable can further be integrated into an active-learning problem with some success [30]. While the added complexity in the observables provides more flexibility in the modeling range of the Koopman operators, it does reduce the effectiveness of active learning. This is a result of more data being required to effectively learn the nonlinear observables. When compared to deep neural network models, the Koopman-based linear model still has a significant advantage in data-efficiency and control through active learning [30].
    V-D Robustness and Stability
    Efforts have also been made to ensure robustness and stability in the face of modeling errors. In [92], the concept of a deep stochastic Koopman operator is introduced to establish stability in nonlinear stochastic systems. This approach involves designing a probabilistic neural network to estimate the distribution of observables, which is then propagated through the Koopman matrices. The training process simultaneously optimizes the parameters of both the probabilistic neural network and the Koopman operator. Further, in [83], researchers compute the nearest stable matrix solution to minimize the least-squares reconstruction error. Rather than addressing uncertainty solely at the modeling stage, alternative approaches explore strategies for considering stability and model mismatch directly during the controller design phase. For instance, modeling errors can be effectively managed by employing a constraint-tightening approach within a proposed tracking MPC strategy [93]. This approach ensures a recursively feasible and input-to-state stable system even in the presence of bounded modeling errors. Another approach to tackle this issue involves constructing a conformant model that combines trace conformance and reach-set conformance [94]. To establish a robust MPC structure, [95] leverages an approximation of the Lipschitz constant for both state- and control-dependent model errors obtained during training. This approach can bound the prediction error across the planning horizon and enable the formulation of an online robust model-based controller.
    VI Applications in Robotic platforms
    In this section, we introduce the challenges involved in modeling different robotic platforms and how Koopman-based implementations can help address them.
    VI-A Aerial Robots
    The inherent nonlinearity of aerial robots presents significant challenges for achieving precise control. Key challenges arise when there is interaction with the environment that is governed by hard-to-model aerodynamic effects [96], such as variable wind gusts and ground effects when landing or flying at a low altitude [97], which can corrupt the nominal model and lead to unexpected performance.
    Data-driven Koopman-based approaches have attempted to mitigate the effect of such uncertain aerodynamic interactions by extracting and adapting online system models used to linearly control the aerial robot. This overall idea can be achieved in different ways. For example, [31] uses episodic learning to estimate the Koopman eigenfunction pairs and obtain the resulting control inputs in real time to handle the ground effect when a multirotor is landing. A sequence of Koopman eigenfunctions is iteratively learned from data. Accordingly, nonlinear control signals are improved from the nominal control law. Another implementation of the Koopman operator theory in the aerial robots is to jointly learn a function dictionary and the lifted Koopman bilinear model to achieve quadrotor trajectory tracking at a low altitude [98]. A neural network is combined with the Koopman operator to update both the lifted states and inputs of the robot with online measurements. In a different effort, a hierarchical structure is proposed that refines the reference signal sent to the high-rate, pre-tuned low-level controller of the aerial robot to deal with uncertainty [53]. A model of the reference and the actual output of the disturbed robot is learned via Koopman operator theory and is utilized in the outer controller to decrease the effect of environmental disturbance in real time.
    VI-B Ground Robots
    Wheeled robots are subject to nonlinear dynamic effects governed by continuous contact with the ground that are hard to model. In many applications, wheeled robots are called to navigate smoothly and efficiently over varying types of terrain, including rugged and deformable terrain [99]. This is usually formulated as an optimization problem over a time horizon, where a model of the underlying system in the changing environment is required. Koopman operator-based methods can serve as the model extraction method without imposing stringent requirements on information to be known (or assumed) a priori. Then, a model-based controller, e.g., Model Predictive Controller (MPC), can be designed based on this data-driven model to achieve a trajectory tracking objective [93]. Other considerations including rotational motions of the mobile wheeled robots are emphasized via coordinate transformation by introducing virtual control inputs [100].
    Legged robots can traverse different and more complicated terrains than wheeled robots. Legged robots usually have highly nonlinear and sophisticated dynamics, while locomotion is governed by discrete impacts, thus making modeling and control a hard task [101] An initial attempt at modeling legged robots with the Koopman operator is done with the quadruped leg dynamics on deformable terrains in [102]. An experimental framework has been proposed to obtain a data-driven Koopman-based model of a quadruped’s leg dynamics as a switched dynamical system. Experimental results have shown that the learned switched system model can be used to predict gait trajectories when operating over unknown terrains. Though current developments in this area are still in their infancy, more research activity is expected, e.g., controller design with Koopman-based models, which may become a potentially powerful tool in legged locomotion.
    Distinctively yet crucially, the Koopman operator has also found its way into the realm of autonomous driving [103]. A key challenge is how to identify a global vehicle dynamics model considering the inherent complexity of the different subsystems and the induced nonlinearities and uncertainties. Considerable efforts have been directed toward utilizing the Koopman operator theory for estimating vehicle models [104]. The resulting approximated models can then be integrated into controller design, such as Model Predictive Control (MPC) [105, 106, 107]. In some cases, like in operation over deformable terrains with significant height variations and the presence of bumps [108], it has been shown possible to obtain a linear representation of both vehicle and terrain interaction dynamics using Koopman estimation. Further advancements have considered bilinear model formations for autonomous vehicles [109] and their combination with deep neural networks [91]. The integration of driver-in-the-loop dynamics has been explored in [110], reflecting a holistic approach to autonomous driving. Additionally, [111] introduces the embedding of the stochastic Koopman operator structure within an attention-based framework. This innovative approach is leveraged for abnormal behavior detection in autonomous driving systems, thereby enhancing the safety of vehicle controllers. The widespread application of the Koopman operator demonstrates its versatility and efficacy in advancing autonomous driving technologies.
    VI-C Robot Manipulators
    (Rigid) robot manipulators are another popular robotic platform because of their stability, feasibility, and safety. Modeling and control for most robot arms have been well established, and these robots are robust to environmental uncertainties. However, recent advances in embodied intelligence consider to a large extent robot manipulators, and even for these systems the simulation-to-real gap remains a challenge. Besides works on model approximation approaches for robot arms [46, 27, 112], Koopman operator theory is usually employed as assistance for other learning methods such as reinforcement learning. For instance, to guide the learning process, Koopman-based methods have been used to build a human intent model from human-demonstrated trajectories [113]. Another example is embedding the Koopman operator into LQR formulation for efficient and flexible policy learning [114].
    VI-D Underwater Robots
    Beyond several works using the Koopman operator to analyze fluid flows directly [115], Koopman-based methods have been also implemented for modeling and control of underwater robotic systems that interact with those changing flows. Underwater robots are typically underactuated and nonlinear because of propulsion designs aimed at reducing weight and cost. On the other hand, such robots often are subject to uncertain disturbance forces and moments that can be generated by fluid fields or other objects [116]. The use of the Koopman operator can introduce linearity in the model expression, which can be further utilized as model constraints in the controller design [117, 118]. Further, feedback policies that can learn or adapt online to unmodeled changes provide directions to solve the aforementioned challenges. In [64], a Koopman-LQR approach is proposed and experimentally tested in the dynamic estimation and control of a tail-actuated robotic fish. Results show that the online-updated data-driven model constructed using the Koopman operator significantly improves control performance in trajectory-tracking tasks compared to a tuned PID controller. Even though Koopman-based methods display better performance than other state-of-the-art data-driven modeling methods, like NARX (neural network approach [119]) and SINDy [120] in this illustration, the control of underwater robots with the Koopman operator is still an open direction in more complicated environments with harder tasks.
    VI-E Soft Robots
    Soft robots have received significant attention in recent years, and have grown into an important research topic. In contrast to their rigid counterparts, soft robots have bodies made out of intrinsically soft or extensible materials, which exhibit unprecedented adaptation to complex environments and can absorb impact energy for safe interactions with the environment, other robots, and even humans. Soft robots are particularly well-suited for data-driven modeling approaches because in most cases they can be operated even under randomized control inputs without posing much of a physical risk to themselves or their surroundings. Data can thus be collected from them much more safely than from traditional rigid-body robots.
    Koopman operator theory provides a data-driven alternative to traditional modeling and control methods for soft robots. It is attractive for soft robot applications because it generates global linear representations of their dynamics (which may be otherwise intractable) that are compatible with traditional control techniques such as MPC. Koopman theory has been applied to the modeling of various soft robotic systems, enabling real-time control [121, 122, 123, 124, 125, 92, 48, 126, 127].
    Major challenges remain in applying Koopman methods to the modeling and control of soft robots. One disadvantage of Koopman-based linearization is that it involves increasing the dimension of the state space to achieve linearity. Soft robots and other continuum systems have inherently high-dimensional dynamics, therefore it is not always possible to linearize their dynamics through lifting without exceeding reasonable computational complexity. Instead, low-dimensional parameterizations of soft robots must be utilized to make them compatible with Koopman modeling techniques, which inadvertently rely on assumptions. In turn, the resulting models may not generalize well, especially in contact-rich applications. How to choose the most appropriate description of a soft robot for compatibility with Koopman methods remains an open area of research.
    VI-F Other Types of Robot Platforms
    Beyond the categories described above, Koopman operator theory can also contribute to advancing the modeling and control of other robot platforms. One example is the case of autonomous robotic excavators, where the excavator works and interacts with the surrounding soil in a highly nonlinear and complex manner [128, 129]. These nonlinear bucket-soil dynamics are learned as a set of linear equations with the Koopman operator. System identification and control applied to rehabilitation robots or limb assistive devices [130, 131] is another nascent area of application. Legged robots and manipulator robots involved in rehabilitation often undergo discrete transitions between contact and non-contact states, leading to a switch in governing equations. The application of Koopman operator theory to these hybrid systems offers the advantage of directly converting a diverse array of heterogeneous systems into a unified global linear model [132]. This characteristic proves particularly beneficial for control purposes in managing the intricacies of these dynamic systems. More instances include but are not limited to model extraction of snake-like robots [133], smarticle ensembles [134], and surgery robots [135].
    VI-G Multi-agent Systems
    In addition to its applications in single-agent systems, Koopman operator theory has proven valuable in the domain of multi-agent systems. In a multi-agent robotic system, multiple robots or agents collaborate to execute tasks within a shared environment. A comprehensive introduction and tutorial on the application of Koopman system estimation and control in multi-agent systems is provided in [136, 137].
    VII Discussion and Conclusion

### v2

@@ V-C1 Optimization-based Methods
Note that equation (27) is solved over dictionary $\mathbf{\Psi}$ which belongs to a parametric family of functions; therefore, the optimization problem is generally non-convex, and finding globally optimal solutions is not guaranteed. More importantly, minimizing the EDMD’s residual error does not necessarily lead to a close-to-invariant subspace. Figure 3 (left) illustrates this point with an example of a non-invariant subspace whose residual error can be made arbitrarily small depending on the choice of basis for the subspace. Therefore, optimization (27) may yield models unsuitable for long-term prediction.
@@
The work in [162] addresses this problem by providing an alternative loss function that captures the quality of the subspace. The notion of temporal forward-backward consistency index (or consistency index for brevity) is defined as 777 This definition is equivalent but different from the definition in [162].
@@
The consistency index has several important properties. (1) $\mathcal{I}_{C}\in[0,1]$; (2) Unlike the cost function in (27), $\mathcal{I}_{C}$ only depends on the subspace $\operatorname{span}(\mathbf{\Psi})$ and not any specific basis (Fig. 3 illustrates this); (3) Under specific change of basis, $\mathcal{I}_{C}$ can be viewed as the maximum eigenvalue of a positive semidefinite matrix ($I-K_{F}K_{B}$ is not generally symmetric) enabling the use of common optimization solvers; (4) Importantly, it provides a tight upper bound on the relative prediction error of EDMD on the entire subspace. Formally,
@@
Figure 3: Left: EDMD’s residual error for linear system $x^{+}=0.6x$ and the family of dictionaries $D_{\beta}(x)=[x,x+\beta\sin(x)]$ with $\beta\in[0.01,100]$. Note that for all $\beta\in\mathbb{R}\setminus\{0\}$, all dictionaries span the same subspace $\mathcal{S}=\operatorname{span}\{x,\sin(x)\}$. The residual error depends on the choice of basis for subspace $\mathcal{S}$. More importantly, $\mathcal{S}$ is not Koopman-invariant but the residual error can be arbitrarily close to zero depending on the basis. Right: The square root of consistency index for the system and family of dictionaries. Unlike EDMD’s residual error, the consistency index does not depend on the choice of basis and accurately measures the approximation quality of the subspace.
@@
which minimizes the maximum EDMD function prediction error over the subspace. Note that, by minimizing the consistency index, one accounts for accurate prediction of all uncountably many members of the function space, as opposed to only finitely many functions in the optimization (27). Therefore, by minimizing the consistency index, one can expect a more accurate prediction, especially in the long term. Figure 4 illustrates this point in a nonlinear pendulum example.
@@
V-C2 Algebraic Search for Koopman-Invariant Subspaces
@@
The work [31] provides a data-driven necessary and almost surely sufficient condition to identify all Koopman eigenfunctions in an arbitrary finite-dimensional function space. This result directly relies on the eigendecomposition of forward and backward EDMD matrices (cf. (V-C1)). Moreover, [31] provides an efficient and provably correct algebraic algorithm termed Symmetric Subspace Decomposition (SSD) that finds the maximal Koopman-invariant subspace in arbitrary finite-dimensional function spaces (termed “search space”). [32] provides a parallel version of the SSD algorithm suitable for searching through high-dimensional spaces. Given as search space the finite-dimensional function space spanned by the dictionary $\mathbf{\Psi}_{s}$, any basis $\mathbf{\Psi}$ with elements in $\operatorname{span}(\mathbf{\Psi}_{s})$ can be described using a matrix $C$ with full column rank as $\mathbf{\Psi}^{T}(\cdot)=\mathbf{\Psi}_{s}^{T}(\cdot)C$. If the subspace $\operatorname{span}(\mathbf{\Psi})$ is invariant under the Koopman operator, this gets reflected in the data, $\mathcal{R}(\mathbf{\Psi}(X)^{T})=\mathcal{R}(\mathbf{\Psi}(Y)^{T})$ (where $\mathcal{R}(\cdot)$ stands for the range space of its argument), which can also be written as
@@
VI Open Directions and Challenges